# Characterizing the Hyperuniformity of Disordered Network Metamaterials

### v1

@@ III.1 Two-Dimensional Networks
Here, we present the ensemble-averaged spectral densities $\tilde{\chi}_{{}_{V}}(k)$ obtained using Eq. 6 on pixelized images of V, D, C, and G networks derived from totally uncorrelated, URL, and disordered stealthy hyperuniform point patterns in $\mathbb{R}^{2}$. To compare two-phase structures generated using many combinations of point patterns and tessellation schemes, we use the specific surface $s$—the interfacial area per unit volume—to make length scales dimensionless, as advocated for in Refs. 27, 35.
III.1.1 Totally Uncorrelated Networks
@@
III.1.2 URL Networks
Figure 3 (a) shows the dimensionless spectral density $\tilde{\chi}_{{}_{V}}(k)s^{2}$ as a function of dimensionless wavenumber $k/(2\pi s)$ for V, D, C, and G networks derived from URL point patterns with $a=0.1$. Each curve (and all those discussed in this subsection) is an ensemble average of 100 such networks derived from configurations with $N=50^{2}$ points. For these URL-derived networks with relatively small perturbations, we find that G networks are nonhyperuniform, while the C, D, and V networks appear to have hyperuniform scaling at intermediate length scales and then nonhyperuniform scaling at large length scales. Each of the C, D, and V network ensembles is effectively hyperuniform with $H=2.3\times 10^{-4},1.7\times 10^{-5}$, and $7.0\times 10^{-6}$, respectively.
@@ III.1.2 URL Networks
III.1.3 Disordered Stealthy Networks
In Fig. 5 (a) we show the dimensionless spectral densities $\tilde{\chi}_{{}_{V}}(k)s^{2}$ as a function of dimensionless wavenumber $k/(2\pi s)$ for V, D, C, and G networks derived from “high-$\chi$” disordered stealthy hyperuniform point patterns with $\chi=0.48$. Each curve in this subsection is an ensemble average of 98 networks derived from configurations with $N=40^{2}$ points. The rank-ordering of hyperuniformity degree in Fig. 5 (a) matches the $a>0.7$ URL network results, i.e., G networks are nonhyperuniform, while C, V, and D networks are effectively hyperuniform (in order of increasing $H$). In numerically generated disordered stealthy point patterns, the value of $S(k)$ in the “stealthy region” ($0<k\leq K$) is not identically zero, but some very small number ($\approx 10^{-22}$ in our configurations). Notably, we find that the shape of the “stealthy region” one expects at small-$k$ is present in each of the effectively hyperuniform spectral densities, but at a much larger ordinate values compared to the progenitor point pattern.
@@ III.1.3 Disordered Stealthy Networks
From these results, and those in Secs. III.1.1 and III.1.2, it is clear that the degree to which a network structure inherits the hyperuniformity of its progenitor point pattern depends both on the tessellation scheme and the translational order across length scales of the progenitor point pattern. In particular, while D networks derived from totally uncorrelated point patterns have the greatest density fluctuations at intermediate and large length scales, this is not the case for the URL and stealthy point patterns because D networks are able to partially inherit the hyperuniformity of those point patterns, while G networks do not. Thus, across networks derived from hyperuniform point patterns, G networks have the greatest large-scale density fluctuations.
@@
Another important trend is evident in Fig. 3 (c), which demonstrates that some tessellation schemes are better at preserving the hyperuniformity of disordered hyperuniform systems with greater degrees of local translational order than others. In particular, V tessellations better preserve the hyperuniformity of disordered hyperuniform systems with greater degrees of local translational order, while C tessellations better preserve the hyperuniformity of those with lesser degrees of local translational order. The sensitivity of the rank-ordering of the degree of hyperuniformity to the tessellation and point pattern type is consistent with changes observed in the ranking-ordering of certain physical properties across networks generated via different tessellations of different point patterns [14].
III.2 Edge-Length Variance $\sigma^{2}_{\ell}(R)$
To ensure that the degree of hyperuniformity of networks derived from hyperuniform URL and disordered stealthy point patterns is not being impacted by mapping these networks to two-phase media, we require a way to characterize the network structure directly. To do so, we propose computing the variance of the total edge length of a network embedded in $\mathbb{R}^{d}$ within a uniformly randomly positioned observation window of radius $R$, which we term the edge-length variance $\sigma^{2}_{\ell}(R)$. Such a characterization method is motivated by the collection of variance measurements that were recently shown to be effective order metrics for point patterns [29], two-phase media [27], and scalar fields [28]. Since one of the two “phases” in such a network system would have a vanishing volume fraction, instead of using $s$ to make distances dimensionless, we use $d-1]{\rho_{\ell}}$, where $\rho_{\ell}$ is the edge length per unit volume, or the edge-length density.
While the spectral density methods used above are reasonably fast—on the order of several minutes per structure for the system sizes used herein—they suffer from potential discretization errors, especially when generating the pixelizations of very large or nearly crystalline networks. Moreover, one must choose a particular beam thickness (and shape for $d>2$) which may impact the density fluctuations of the structure in unexpected ways. Both of the aforementioned issues can be avoided when computing $\sigma^{2}_{\ell}(R)$ because the intersections between line segments (the network edges) and hyperspheres (the observation window of radius $R$) can be computed exactly with relative ease and no decoration of the network structure is required to do the computations. The volume fraction variance can also be computed without discretizing the structure, but due to the increased complexity of the geometries of two-phase media, the computation is much more computationally intensive than the corresponding one for $\sigma^{2}_{\ell}(R)$, where computing $\sigma^{2}_{\ell}(R)$ for a structure derived from a $N=300^{2}$ 2D point pattern takes on the order of an hour while the corresponding volume fraction variance calculation takes several hours.
@@ III.2 Edge-Length Variance $\sigma^{2}_{\ell}(R)$
In Figs. 6 (a) and (b), we show the $\sigma^{2}_{\ell}(R)$ curves for 50-configuration ensembles of V, D, C, and G networks derived from totally uncorrelated point patterns with $N=300^{2}$ and 50-configuration ensembles of V, D, C, and G networks derived from URL configurations with $a=0.5$ and $N=200^{2}$. In Fig. 6 (c) the curves correspond to 98-configuration ensembles of V, D, C, and G networks derived from disordered stealthy hyperuniform point patterns with $\chi=0.48$ and $N=100^{2}$. In each subfigure of Fig. 6, the small-$R$ behavior of each set of curve collapses, indicating that $d-1]{\rho_{\ell}}$ is indeed a reasonable choice of scale. From each of the four curves in Fig. 6 (a) and the G curves in Figs. 6 (b) and (c), it is evident that the nonhyperuniform systems examined in Sec. III.1 have $\sigma^{2}_{\ell}(R)\sim R^{2}$ (i.e., scale like the observation window volume) at large $R$. Like in Fig. 2, one can clearly observe that the ordering of the totally uncorrelated point pattern networks from largest to smallest large-scale density fluctuations is D $>$ G $>$ V $>$ C. Moreover, the V, D, and C curves in Figs. 6 (b) and (c) exhibit $\sigma^{2}_{\ell}(R)\sim R$ scaling at intermediate length scales, which increases at larger length scales. This change in scaling behavior is consistent with the onset of nonhyperuniform scaling at small $k$ evident in Figs. 3 (b) and 5 (a).
From these results, it is clear that $\sigma^{2}_{\ell}(R)$ follows the same scaling behaviors as the number variance $\sigma_{N}^{2}(R)$, i.e., nonhyperuniform systems have $R^{d}$ scaling while hyperuniform systems have slower scaling. These results also indicate that the onset of nonhyperuniformity observed at large length scales in Sec. III.1 is not due to mapping these networks to two-phase media, but intrinsic to the network structures themselves. Importantly, this implies that the particular choice of parameters used to convert the network to a two-phase medium does not significantly impact the hyperuniformity of the resulting structure and one can probe the network structure directly to assess the hyperuniformity of structures derived from it.
III.3 Three-Dimensional Networks
Having established that $\sigma^{2}_{\ell}(R)$ faithfully captures the intermediate and large-scale density fluctuations in 2D network structures, we now apply this same method to 3D networks. We opt not to map the 3D networks to two-phase media and characterize the resulting structures with either the volume-fraction variance or spectral density due to the significantly increased difficulty of exactly describing the geometry of the media at the network vertices, where several of these beams overlap compared to $d=2$. Calculations involving the newly proposed edge-length variance are much easier to execute, by comparison, for the reasons stated in the previous subsection.
@@ III.3 Three-Dimensional Networks
In Fig. 7 (a) we present the edge-length variance scaled by the window volume $\sigma_{\ell}^{2}(R)/R^{3}$ as a function of the scaled window radius $R\rho_{\ell}^{1/2}$ for 500-configuration ensembles of V, D, C, and G networks derived from 3D totally uncorrelated point patterns with $N=40^{3}$ points. Here, we scale $\sigma_{\ell}^{2}(R)$ by the observation window volume $R^{3}$ for the sake of visual clarity. Curves that plateau at large $R$ correspond to nonhyperuniform structures and those that decrease at large $R$ correspond to hyperuniform structures. Figure 7 (a) reveals that the the networks derived from totally uncorrelated point patterns have nonhyperuniform $\sigma_{\ell}^{2}(R)\sim R^{3}$ scaling consistent the the nonhyperuniformity of the progenitor point pattern. However, the rank-ordering of large-scale density fluctuations in these totally uncorrelated networks changes from D $>$ G $>$ V $>$ C for the $d=2$ case (cf. Fig. 6 (a)), to D $>$ G $>$ C $>$ V.
@@
Figures 7 (b), and (c) show the edge-length variance scaled by the window volume $\sigma_{\ell}^{2}(R)/R^{3}$ as a function of the scaled window radius $\sigma^{2}_{\ell}(R)$ for 500-configuration ensembles of V, D, C, and G networks derived from 3D URL point patterns with $N=40^{3}$ points and $a=0.5$ and $1.0$, respectively. In these figures, we observe the expected behavior of the G networks, i.e., they have more obvious nonhyperuniform scaling than the V, C, and D networks and the greatest density fluctuations at large $R$. The rank-ordering of the other three networks types for stealthy and large-$a$ URL point patterns, however, is different from the 2D case. Specifically, while D is still the most weakly hyperuniform of the three, V networks are consistently more hyperuniform than the corresponding C networks at large $R$. The change observed here in the rank-ordering between $d=2$ and $d=3$ is not surprising considering that the structures that minimize large-scale density fluctuations in point patterns also change across dimensions (e.g., the $A_{2}$ lattice minimizes such fluctuations in $d=2$, while $A_{3}^{*}$ is conjectured to do so in $d=3$ [18]). We also note that, on intermediate length scales, D networks appear to have the greatest density fluctuations of the four network types, regardless of the (non)hyperuniformity of the progenitor point pattern. In addition, while the onset of nonhyperuniform scaling at large length scales is visible in 7 (b), it is less obvious in 7 (c), suggesting that the transition to nonhyperuniform scaling occurs at larger length scales in systems with greater degrees of disorder on small length scales.
IV Discussion and Conclusions
@@
In this work, we probed the ability of the first step of the configuration-to-network-to-3D printed structure pipeline devised by Obrero et al. [17] to generate hyperuniform network structures. In particular, we generated Voronoi, Delaunay, Delaunay-Centroidal, and Gabriel tessellations of totally uncorrelated, uniformly randomized lattice, and disordered stealthy hyperuniform point patterns in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. We then characterized the structures of the networks generated by the connected sets of edges in the aforementioned tessellations in two ways. First, for networks in $\mathbb{R}^{2}$, we created two-phase media by mapping each edge in the network to a rectangle and then computed the spectral density $\tilde{\chi}_{{}_{V}}(k)$ of the resulting structure. We then proposed a new variance-based measure inspired by other recent variance-based order metrics [27, 28, 29], the edge-length variance $\sigma_{\ell}^{2}(R)$, to characterize the structures of spatially embedded networks directly without needing to convert to structure to a two-phase medium. We also suggested the use of the $(d-1)$th root of the edge-length density $\rho_{\ell}$ to make length scales dimensionless when comparing different types of networks. We then applied this new characterization method to the networks derived from the aforementioned point patterns in both $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$.
@@
By examining the spectral densities of two-phase media derived from 2D networks, we found that networks derived from nonhyperuniform point patterns inherit the nonhyperuniformity of the progenitor point pattern, but tessellations of hyperuniform point patterns only partially inherit the hyperuniformity of the progenitor point pattern (or not at all, in the case of G tessellations). In particular, we found that C, V, and D tessellations have hyperuniform density fluctuations at intermediate length scales, but nonhyperuniform density fluctuations on sufficiently large length scales. Therefore, these network structures are, at best, effectively hyperuniform and in other cases nonhyperuniform. For disordered hyperuniform systems with a large degree of local translational order, V networks best preserve hyperuniformity, while C networks best preserve the hyperuniformity of disordered hyperuniform point patterns with a small degree of local translational order. We also found that, despite hyperuniformity describing the large-scale structure of a system, a network derived from a hyperuniform point pattern with a small degree of local translational order will be more weakly hyperuniform than one derived from a point pattern that has the same strength of hyperuniformity but has a greater degree of local translational order. Moreover, the rank-ordering of the degree of hyperuniformity in network structures is sensitive to the particular tessellation scheme and both the small- and large-scale translational order of the progenitor point pattern. Such a sensitivity is consistent with the sensitivity of the rank-ordering of the transport and elastic properties of these network structures to the same variables [14].
@@
We then showed that the edge-length variance $\sigma_{\ell}^{2}(R)$ faithfully captures the intermediate- and large-scale structural characteristics of the networks when treated as a set of line segments (i.e., not treated as a two-phase medium) and has the same $R^{d}$ scaling for nonhyperuniform systems and slower scaling for hyperuniform systems as the number variance $\sigma_{N}^{2}(R)$. The calculation of $\sigma_{\ell}^{2}(R)$ is much faster and simpler compared to the corresponding spectral density or volume-fraction variance calculations due to relative simplicity of the line-segment-based network structure. Moreover, we showed that the loss of hyperuniformity at large length scales observed in the spectral densities of V, D, and C networks derived from 2D hyperuniform point patterns is also present in the $\sigma_{\ell}^{2}(R)$ curves, suggesting that this lack of hyperuniformity is intrinsic to the networks structures. This finding suggests that the particular choice of beam shape when designing two-phase structures for manufacturing purposes will not significantly impact the large-scale density fluctuations of the manufactured structure—an important practical consideration. We then applied $\sigma_{\ell}^{2}(R)$ to 3D networks derived from nonhyperuniform and disordered hyperuniform point patterns and found that V networks tend to be the best at suppressing large-scale density fluctuations. This change in optimal structure from C or V networks for $d=2$ to $V$ for $d=3$ is not unexpected given that the currently known structures that minimize large-scale density fluctuations in point patterns also change across space dimensions [18].
@@
The results presented here can be used to inform the design of effectively hyperuniform network metamaterials in two and three space dimensions for future theoretical and experimental characterization of their physical properties, e.g., their thermal conductivity, electrical conductivity, and acoustic properties. Moreover, our results demonstrating the nonhyperuniformity of networks derived from the spatial tessellations of hyperuniform point patterns on large length scales motivates a search for network structures that are truly hyperuniform as opposed to effectively hyperuniform. Possible inverse-design routes for this hyperuniform network structure search include optimization schemes (e.g., simulated annealing) or machine learning methods. Another important area of future work is the development of efficient methods to construct and characterize the two-phase media associated with 3D networks to better study the structure and properties of 3D-printed network metamaterials.

### v2

@@ IV.1 Two-dimensional networks
Here, we present the ensemble-averaged spectral densities $\tilde{\chi}_{{}_{V}}(k)$ obtained using Eq. 6 on pixelized images of V, D, C, and G networks derived from totally uncorrelated, URL, and disordered stealthy hyperuniform point patterns in $\mathbb{R}^{2}$. To compare two-phase structures generated using many combinations of point patterns and tessellation schemes, we use the specific surface $s$—the interfacial area per unit volume—to make length scales dimensionless, as advocated for in Refs. 31, 41.
IV.1.1 Totally uncorrelated networks
@@
IV.1.2 URL networks
Figure 3 (a) shows the dimensionless spectral density $\tilde{\chi}_{{}_{V}}(k)s^{2}$ as a function of dimensionless wavenumber $k/(2\pi s)$ for V, D, C, and G networks derived from URL point patterns with $a=0.1$. Each curve (and all those discussed in this subsection) is an ensemble average of 100 such networks derived from configurations with $N=50^{2}$ points. For these URL-derived networks with relatively small perturbations, we find that G networks are nonhyperuniform, while the C, D, and V networks are effectively hyperuniform. The hyperuniformity index $H$ for each of the C, D, and V network ensembles is $2.3\times 10^{-4},1.7\times 10^{-5}$, and $7.0\times 10^{-6}$, respectively.
@@ IV.1.2 URL networks
IV.1.3 Disordered stealthy networks
In Fig. 5 (a) we show the dimensionless spectral densities $\tilde{\chi}_{{}_{V}}(k)s^{2}$ as a function of dimensionless wavenumber $k/(2\pi s)$ for V, D, C, and G networks derived from “high-$\chi$” disordered stealthy hyperuniform point patterns with $\chi=0.48$. Each curve in this subsection is an ensemble average of 98 networks derived from configurations with $N=40^{2}$ points. The rank-ordering of hyperuniformity degree in Fig. 5 (a) matches the $a>0.7$ URL network results, i.e., G networks are nonhyperuniform, while C, V, and D networks are effectively hyperuniform (in order of increasing $H$). In numerically generated disordered stealthy point patterns, the value of $S(k)$ in the “stealthy region” ($0<k\leq K$) is not identically zero, but some very small number ($\approx 10^{-22}$ in our configurations). Notably, we find that the shape of the “stealthy region” one expects at small-$k$ is present in each of the effectively hyperuniform spectral densities, but at much larger ordinate values compared to the progenitor point pattern.
@@ IV.1.3 Disordered stealthy networks
From these results, and those in Secs. IV.1.1 and IV.1.2, it is clear that the degree to which a network structure inherits the hyperuniformity of its progenitor point pattern depends both on the tessellation scheme and the translational order across length scales of the progenitor point pattern. In particular, while D networks derived from totally uncorrelated point patterns have the greatest density fluctuations at intermediate and large length scales, this is not the case for the URL and stealthy point patterns because D networks are able to partially inherit the hyperuniformity of those point patterns, while G networks do not. Thus, across networks derived from hyperuniform point patterns, G networks have the greatest large-scale density fluctuations.
@@
Another important trend is evident in Fig. 3 (c), which demonstrates that some tessellation schemes are better at preserving the hyperuniformity of disordered hyperuniform systems with greater degrees of local translational order than others. In particular, V tessellations better preserve the hyperuniformity of disordered hyperuniform systems with greater degrees of local translational order, while C tessellations better preserve the hyperuniformity of those with lesser degrees of local translational order. The sensitivity of the rank-ordering of the degree of hyperuniformity to the tessellation and point pattern type is consistent with changes observed in the rank-ordering of certain physical properties across networks generated via different tessellations of different point patterns [14].
IV.2 Edge-length variance $\sigma^{2}_{\ell}(R)$
Here, we apply the edge-length variance $\sigma_{\ell}^{2}(R)$ to our spatial network structures, to probe their density fluctuations without converting them into two-phase media. In analogy with the use of the specific surface $s$ to scale distances in two-phase media, we use $d-1]{\rho_{\ell}}$, where $\rho_{\ell}$ is the edge length per unit volume, or the edge-length density, to make distances dimensionless in the analysis of our networks. Moreover, we note that we use a maximum window radius of $L/4$, where $L$ is the side length of the square (cubic) simulation box, to mitigate finite-size effects.
@@ IV.2 Edge-length variance $\sigma^{2}_{\ell}(R)$
In Figs. 6 (a) and (b), we show the $\sigma^{2}_{\ell}(R)$ curves for 50-configuration ensembles of V, D, C, and G networks derived from totally uncorrelated point patterns with $N=300^{2}$ and 50-configuration ensembles of V, D, C, and G networks derived from URL configurations with $a=0.5$ and $N=200^{2}$. In Fig. 6 (c) the curves correspond to 98-configuration ensembles of V, D, C, and G networks derived from disordered stealthy hyperuniform point patterns with $\chi=0.48$ and $N=100^{2}$. In each subfigure of Fig. 6, the small-$R$ behavior of each set of curves collapses, indicating that $d-1]{\rho_{\ell}}$ is indeed a reasonable choice of scale. From each of the four curves in Fig. 6 (a) and the G curves in Figs. 6 (b) and (c), it is evident that the nonhyperuniform systems examined in Sec. IV.1 have $\sigma^{2}_{\ell}(R)\sim R^{2}$ (i.e., scale like the observation window volume) at large $R$. Like in Fig. 2, one can clearly observe that the ordering of the totally uncorrelated point pattern networks from largest to smallest large-scale density fluctuations is D $>$ G $>$ V $>$ C. Moreover, the V, D, and C curves in Figs. 6 (b) and (c) exhibit $\sigma^{2}_{\ell}(R)\sim R$ scaling at intermediate length scales, which increases at larger length scales. This change in scaling behavior is consistent with the onset of degraded hyperuniform scaling at small $k$ evident in Figs. 3 (b) and 5 (a).
From these results, we have confirmed the scaling behavior of $\sigma^{2}_{\ell}(R)$ given in Eq. (15) that matches the number variance $\sigma_{N}^{2}(R)$ scalings, i.e., nonhyperuniform systems have $R^{d}$ scaling while hyperuniform systems have slower scaling. These results also indicate that the onset of the degradation of hyperuniformity observed at large length scales in Sec. IV.1 is not due to mapping these networks to two-phase media, but intrinsic to the network structures themselves. Importantly, this implies that the particular choice of parameters used to convert the network to a two-phase medium does not significantly impact the hyperuniformity of the resulting structure and one can probe the network structure directly to assess the hyperuniformity of structures derived from it.
IV.3 Three-dimensional networks
Having established that $\sigma^{2}_{\ell}(R)$ faithfully captures the intermediate and large-scale density fluctuations in 2D network structures, we now apply this same method to 3D networks. We opt not to map the 3D networks to two-phase media and characterize the resulting structures with either the volume-fraction variance or spectral density due to the significantly increased difficulty of exactly describing the geometry of the media at the network vertices, where several of these beams overlap compared to $d=2$. Calculations involving the newly proposed edge-length variance are much easier to execute, by comparison, for the reasons stated in Sec. III.
@@ IV.3 Three-dimensional networks
In Fig. 7 (a) we present the edge-length variance scaled by the window volume $\sigma_{\ell}^{2}(R)/R^{3}$ as a function of the scaled window radius $R\rho_{\ell}^{1/2}$ for 500-configuration ensembles of V, D, C, and G networks derived from 3D totally uncorrelated point patterns with $N=40^{3}$ points. Here, we scale $\sigma_{\ell}^{2}(R)$ by the observation window volume $R^{3}$ for the sake of visual clarity. Curves that plateau at large $R$ correspond to nonhyperuniform structures and those that decrease at large $R$ correspond to hyperuniform structures. Figure 7 (a) reveals that the networks derived from totally uncorrelated point patterns have nonhyperuniform $\sigma_{\ell}^{2}(R)\sim R^{3}$ scaling consistent with the nonhyperuniformity of the progenitor point pattern. However, the rank-ordering of large-scale density fluctuations in these totally uncorrelated networks changes from D $>$ G $>$ V $>$ C for the $d=2$ case (cf. Fig. 6 (a)), to D $>$ G $>$ C $>$ V.
@@
Figures 7 (b), and (c) show the edge-length variance scaled by the window volume $\sigma_{\ell}^{2}(R)/R^{3}$ as a function of the scaled window radius $\sigma^{2}_{\ell}(R)$ for 500-configuration ensembles of V, D, C, and G networks derived from 3D URL point patterns with $N=40^{3}$ points and $a=0.5$ and $1.0$, respectively. In these figures, we observe the expected behavior of the G networks, i.e., they have more obvious nonhyperuniform scaling than the V, C, and D networks and the greatest density fluctuations at large $R$. The rank-ordering of the other three network types for stealthy and large-$a$ URL point patterns, however, is different from the 2D case. Specifically, while D is still the most weakly hyperuniform of the three, V networks are consistently more hyperuniform than the corresponding C networks at large $R$. The change observed here in the rank-ordering between $d=2$ and $d=3$ is not surprising considering that the structures that minimize large-scale density fluctuations in point patterns also change across dimensions (e.g., the $A_{2}$ lattice minimizes such fluctuations in $d=2$, while $A_{3}^{*}$ is conjectured to do so in $d=3$ [18]). We also note that, on intermediate length scales, D networks appear to have the greatest density fluctuations of the four network types, regardless of the (non)hyperuniformity of the progenitor point pattern. In addition, while the onset of the degraded hyperuniform scaling at large length scales is visible in 7 (b), it is less obvious in 7 (c), suggesting that the degradation of hyperuniformity occurs at larger length scales in systems with greater degrees of disorder on small length scales.
V Discussion and Conclusions
@@
In this work, we probed the ability of the first step of the configuration-to-network-to-3D printed structure pipeline devised by Obrero et al. [17] to generate hyperuniform network structures. In particular, we generated Voronoi, Delaunay, Delaunay-Centroidal, and Gabriel tessellations of totally uncorrelated, uniformly randomized lattice, and disordered stealthy hyperuniform point patterns in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. We then characterized the structures of the networks generated by the connected sets of edges in the aforementioned tessellations in two ways. First, for networks in $\mathbb{R}^{2}$, we created two-phase media by mapping each edge in the network to a rectangle and then computed the spectral density $\tilde{\chi}_{{}_{V}}(k)$ of the resulting structure. We then proposed a new variance-based measure inspired by other recent variance-based characterizations of disorder across length scales [31, 32, 33], the edge-length variance $\sigma_{\ell}^{2}(R)$, to characterize the structures of spatially embedded networks directly without needing to convert to structure to a two-phase medium. We also suggested the use of the $(d-1)$th root of the edge-length density $\rho_{\ell}$ to make length scales dimensionless when comparing different types of networks. We then applied this new characterization method to the networks derived from the aforementioned point patterns in both $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$.
@@
By examining the spectral densities of two-phase media derived from 2D networks, we found that networks derived from nonhyperuniform point patterns inherit the nonhyperuniformity of the progenitor point pattern, but tessellations of hyperuniform point patterns only partially inherit the hyperuniformity of the progenitor point pattern (or not at all, in the case of G tessellations). We found that the C, V, and D network structures are, at best, effectively hyperuniform and in other cases nonhyperuniform. For disordered hyperuniform systems with a large degree of local translational order, V networks best preserve hyperuniformity, while C networks best preserve the hyperuniformity of disordered hyperuniform point patterns with a small degree of local translational order. We also found that, despite hyperuniformity describing the large-scale structure of a system, a network derived from a hyperuniform point pattern with a small degree of local translational order will be more weakly hyperuniform than one derived from a point pattern that has the same strength of hyperuniformity but has a greater degree of local translational order. Moreover, the rank-ordering of the degree of hyperuniformity in network structures is sensitive to the particular tessellation scheme and both the small- and large-scale translational order of the progenitor point pattern. Such a sensitivity is consistent with the sensitivity of the rank-ordering of the transport and elastic properties of these network structures to the same variables [14].
@@
We then showed that the edge-length variance $\sigma_{\ell}^{2}(R)$ faithfully captures the intermediate- and large-scale structural characteristics of the networks when treated as a set of line segments (i.e., not treated as a two-phase medium) and has the same $R^{d}$ scaling for nonhyperuniform systems and slower scaling for hyperuniform systems as the number variance $\sigma_{N}^{2}(R)$. The calculation of $\sigma_{\ell}^{2}(R)$ is much faster and simpler compared to the corresponding spectral density or volume-fraction variance calculations due to the relative simplicity of the line-segment-based network structure. Moreover, we showed that the degradation of hyperuniformity at large length scales observed in the spectral densities of V, D, and C networks derived from 2D hyperuniform point patterns is also present in the $\sigma_{\ell}^{2}(R)$ curves, suggesting that this degradation of hyperuniformity is intrinsic to the networks structures. This finding suggests that the particular choice of beam shape when designing two-phase structures for manufacturing purposes will not significantly impact the large-scale density fluctuations of the manufactured structure—an important practical consideration. We then applied $\sigma_{\ell}^{2}(R)$ to 3D networks derived from nonhyperuniform and disordered hyperuniform point patterns and found that V networks tend to be the best at suppressing large-scale density fluctuations. This change in optimal structure from C or V networks for $d=2$ to $V$ for $d=3$ is not unexpected given that the currently known structures that minimize large-scale density fluctuations in point patterns also change across space dimensions [18].
We note here that, while we have only run numerical experiments to examine the phenomena discussed in this paper, we are reasonably confident that the behaviors we observe in the large-scale density fluctuations of our network structures are not numerical artifacts. However, this lack of theoretical background motivates the study of the link between some spectral function and the edge-length variance (like the link between the spectral density and the volume fraction variance). Such a link would be of great interest to both the metamaterial and network science communities.
@@
The results presented here can be used to inform the design of effectively hyperuniform network metamaterials in two and three space dimensions for future theoretical and experimental characterization of their physical properties, e.g., their thermal conductivity, electrical conductivity, and acoustic properties. Another potentially interesting area of future research is the application of $\sigma_{\ell}^{2}(R)$ to existing network structures that have been shown to possess desirable physical properties, e.g., the disordered 2D networks with hyperuniform vertex locations described in Refs. 25 and 26, or the naturally occurring looped leaf vein structures discussed in Ref. [42]. Moreover, our results demonstrating the effective hyperuniformity or nonhyperuniformity of networks derived from the spatial tessellations of various hyperuniform point patterns on large length scales motivate a search for network structures that are truly hyperuniform as opposed to effectively hyperuniform. Possible inverse-design routes for this hyperuniform network structure search include optimization schemes (e.g., simulated annealing) or machine learning methods. Another important area of future work is the development of efficient methods to construct and characterize the two-phase media associated with 3D networks to better study the structure and properties of 3D-printed network metamaterials.